# Byzantine Reliable Broadcast with Low Communication and Time Complexity

### v1

@@ Abstract.
This paper introduces a novel mechanism to lower the communication and computational complexity. Two algorithms are presented that employ this mechanism to reliably broadcast messages in an asynchronous network where less than a third of all nodes are Byzantine. The first algorithm reduces the overhead factor to $2$ and has a time complexity of $3$ if the sender is honest, whereas the second algorithm attains an optimal time complexity of $2$ with the same overhead factor in the absence of equivocation. Moreover, an optimization for real-world implementations is proposed, reducing the overhead factor to $3/2$ under normal operation. Lastly, a lower bound is proved that an overhead factor lower than $3/2$ cannot be achieved for a relevant class of reliable broadcast algorithms.
@@ 1. Introduction
Curiously, there is little work on minimizing the constant in the $O(|m|n)$ term, although it likely dominates the actual bandwidth consumption in practice and is therefore crucial for real-world applications. The design by Cachin and Tessaro is based on a $(n,t+1)$-erasure code, where $t<n/3$ is the largest number of nodes that may exhibit Byzantine behavior. In this design, nodes broadcast encoded fragments of size $|m|/(t+1)\approx 3|m|/n$ instead of broadcasting $m$, sending at least $3|m|n$ bits altogether. An overhead factor of $3$—compared to the ideal scenario where each node receives exactly $|m|$ bits—may already prove too costly for applications with strict bandwidth constraints. All schemes that follow this approach naturally inherit the overhead inherent in this design.
@@
As an added benefit, the transition from a $(n,t+1)$-erasure code to a $(n,2t+1)$-erasure code results in substantial performance improvements, both with respect to the encoding and decoding of fragments. Figure 1 shows performance numbers when encoding/decoding a message of size 1MiB for various network sizes and the two different parameterizations. The experiments were conducted on standard hardware using an optimized library [^1] See https://github.com/AndersTrier/reed-solomon-simd. for Reed-Solomon codes (Reed and Solomon, 1960). The figure confirms the practicality of erasure coding due to its efficiency, exhibiting encoding and decoding times in the millisecond range, and scalability with respect to the network size. Moreover, the figure reveals that the encoding (decoding) time is lower by a factor of 2.2-2.4 (1.8-2.0) for all network sizes when using an $(n,2t+1)$-erasure code. Another interesting property is that the variance is also significantly lower, leading to more predictable encoding and decoding times.
Figure 1. Given an input of size 1MiB, encoding and decoding times for erasure codes with parameters $(n,t+1)$ and $(n,2t+1)$ are shown for different network sizes. The circles and bars mark the averages and the 5th and 95th percentiles, respectively, of the measured encoding and decoding times over 1000 runs.
The paper is organized as follows. The model is formally introduced in §2. A concrete algorithm that makes use of the novel mechanism is presented in §3. Its communication complexity is $2|m|n$ plus a term independent of $|m|$. Given an honest (i.e., non-faulty) sender, each honest node obtains the message after $3$ rounds of communication. A natural question is whether there is an algorithm with an overhead factor lower than $3$ and that terminates in an optimal $2$ rounds given an honest sender. In §4, this question is answered in the affirmative by presenting an algorithm with this property and the same overhead factor if the sender does not equivocate. Otherwise, the worst-case overhead factor is $\frac{5}{2}$ with respect to the largest fragment size transmitted by honest nodes. While the first algorithm only requires collision-resistant hash functions, the second algorithm makes use of threshold signatures. Another aspect of practical relevance is the (worst-case) storage requirement at each node. Both algorithms bound the required space to a small multiple of a predetermined maximum message size. Practical considerations with the goal of reducing the communication complexity further in real-world deployments and simplifying the implementation are discussed in §5. A key observation is that the overhead can be reduced to $\frac{3}{2}$ during periods of synchrony and in the absence of failures. In §6, a lower bound is presented, showing that an overhead factor of $\frac{3}{2}$ is optimal for a particular class of algorithms. Related work is summarized in §7 before the paper concludes in §8.
@@ 2. Model
  • Validity: If the sender is honest and broadcasts message $m$, then every honest node eventually delivers $m$.
@@
  • Totality: If an honest node delivers message $m$, then all honest nodes eventually deliver $m$.
@@
In general, nodes may engage in multiple reliable broadcasts in parallel. In this case, the integrity condition effectively requires that messages are attributable to a uniquely identifiable execution context. To this end, we define that every reliable broadcast execution has a unique identifier. In the following, we assume that every message implicitly contains an identifier and that each node runs a separate instance of the protocol for each identifier. The sender for a particular execution context is assumed to be globally known.
@@
As mentioned above, the objective is to reliably broadcast some message $m$, where the notation $|m|$ is used to denote the size of the message in bits. Instead of sending $m$, the algorithms introduced in §3 and §4 send fragments of $m$, which can be either chunks of $m$ itself or encoded data derived from $m$. Without loss of generality, we assume that there is some upper bound $\ell_{max}$ on the message size, which in turn limits the fragment size. State-of-the-art algorithms use either erasure codes, which can handle missing data, or error-correcting codes, which support the correction of erroneous data. The algorithms in this paper exclusively use an $(n,k)$-erasure codes with optimal reception efficiency, i.e., exactly $k$ out of $n$ symbols are sufficient to reconstruct a message of $k$ symbols. Throughout this paper, the parameters $n$ and $k$ correspond to the total number of nodes and the number of honest nodes, i.e., $k\coloneqq 2t+1$. We assume that each node has access to the routine get_fragments that takes a message $m$ as input and returns a list of $n$ fragments of size $\frac{|m|}{k}=\frac{|m|}{2f+1}<\frac{3}{2}\frac{|m|}{n}$ each. Furthermore, the nodes use the routine recover_message to recover $m$ given any subset $F$ of the $n$ fragments of size $|F|\geq k=2t+1$. This routine is assumed to always return some message, which may simply be a bit string of zeroes in case of an error, e.g., when the input contains fragments of different sizes.
@@
In addition to fragments, the nodes also send hashes, identifiers, and, in the case of the algorithm presented in §4, also signatures. The hashes are assumed to be cryptographically strong in the sense that it is infeasible to find hash collisions (except with negligible probability). Naturally, it is also assumed that it is computationally infeasible to spoof signatures. Since the cryptographic properties only hold for hashes and signatures of a certain minimum size, we introduce a security parameter $\kappa$ and define that the size of all data types other than fragments is bounded by $O(\kappa)$.
@@
Reliable broadcast algorithms are evaluated against multiple complexity measures. For a specific number $n$ of nodes, number $t$ of Byzantine nodes, and message size $\ell$, the communication complexity $\mathcal{C}(n,t,\ell)$ of an algorithm is the total number of bits sent by all honest nodes in the worst case. In the case of a Byzantine sender that sends fragments of different sizes, we define that $\ell$ is the message size corresponding to the largest fragment sent by an honest node. Given that at least $n\ell$ bits need to be transferred and we are concerned with the overhead for large messages, the primary goal is to minimize $\mathcal{L}(n,t)\coloneqq\lim_{\ell\rightarrow\infty}\frac{\mathcal{C}(n,t,\ell)}{\ell n}$. Additionally, the time complexity of an algorithm is of great practical importance, measuring the duration of an execution when normalizing the maximum message delay to $1$ time unit. According to Definition 2.1, an execution with a Byzantine sender may never terminate. Therefore, we restrict our attention to the case where the sender is honest. Due to the validity condition, this good-case time complexity must be bounded. Lastly, the space complexity is considered as well, which is defined as the number of bits that any honest node stores during the execution of the algorithm in the worst case.
@@ 3.2. Description
Each node executing $\mathcal{A}_{bit}$ locally maintains the data structures $F$, $R$, $H$, and $P$ per execution context. Let $\mathcal{F}$, $\Pi$, and $\mathcal{H}$ denote the set of all possible fragments, Merkle proofs, and hashes, respectively. The hash map $F:\mathcal{H}\times V\rightarrow\mathcal{F}\times\Pi$ stores the fragment $f_{j}\in\mathcal{F}$ and Merkle proof $\pi_{j}\in\Pi$ of node $v_{j}$ for the message with the root hash $h\in\mathcal{H}$. If this fragment is locally available, then $F(h,v_{j})=(f_{j},\pi_{i})$, and $F(h,v_{j})=\bot$ otherwise. This hash map is used to collect fragments with the goal of eventually recovering the corresponding message. The hash map $R:\mathcal{H}\rightarrow 2^{V}$ stores the nodes from which a fragment for root hash $h$ has been obtained. This hash map is needed to determine which nodes may still be missing their fragments once a node is able to recover the message. Conversely, for any node $v$, the hash map $H:V\rightarrow 2^{\mathcal{H}}$ contains the set of hashes for which a proposal or fragment was received from node $v$. The purpose of this data structure is to bound both the communication and space complexity. Lastly, $P:\mathcal{H}\rightarrow 2^{V}$ indicates which nodes have proposed the delivery of the message associated with the root hash $h$. The collected proposals are used to ensure that only fragments with sufficient support are broadcast. All hash maps are initially empty. In addition to the hash maps, each node further uses the Boolean variable $done$, initially $false$, to capture the information whether the execution has terminated, either with or without the delivery of a message. In the latter case, the totality condition implies that no honest node will deliver a message for this execution. Let further $\mathbf{H}\coloneqq\cup_{v\in V}H(v)$ and $h_{max}\coloneqq\arg\max_{h\in\mathbf{H}}|P(h)|$, breaking ties arbitrarily if no single hash received the most proposals.
@@
Algorithm 1 Algorithm $\mathcal{A}_{bit}$: Triggered actions at node $v_{i}$. Initially, $F=R=H=P=\{\}$.
@@
Algorithm 2 describes the state-based actions of algorithm $\mathcal{A}_{bit}$. An honest node $v_{i}$ broadcasts its fragment $f_{i}$, at most once, after collecting $12t+$ 1 proposals for the corresponding Merkle root hash $h$. By contrast, a node $v_{i}$ broadcasts a proposal for a specific root hash $h$ not only when it receives its fragment from $sender$ but also when at least $t+1$ fragments for the root hash $h=h_{max}$ have been received and $v_{i}$ has not broadcast the proposal before. Lastly, if a node $v_{i}$ receives at least $12t+$ 1 fragments and proposals for $h_{max}$, it recovers the message $m$ and, additionally, recomputes the fragments and the corresponding root hash. If the computed hash matches $h_{max}$, node $v_{i}$ concludes that the fragments are valid and the message $m$ can be delivered. In this case, $v_{i}$ first recomputes the Merkle proof $\pi_{j}$ and sends $fragment(h_{max},j,f_{j},\pi_{j})$ to each $v_{j}\in V\setminus R(h_{max})$ before calling deliver($m$) to deliver the message. Note that $done$ is set to $true$ after executing these steps even if $h\neq h_{max}$ because it is computationally infeasible to find fragments such that equality holds for some subsets of fragments of size $12t+$ 1.
@@
Algorithm 2 Algorithm $\mathcal{A}_{bit}$: State-based actions at node $v_{i}$. Initially, $F=R=P=\{\}$, $done=false$. Let $h_{max}\coloneqq\arg\max_{h\in\mathbf{H}}|P(h)|$.
@@ 3.3. Analysis
In this section, we prove the correctness of $\mathcal{A}_{bit}$ as well as its communication, time, and space complexity. A series of lemmas is used to simplify the proof structure. The first lemma is concerned with the fragments that honest nodes broadcast, showing that the proposal mechanism successfully ensures that honest nodes only broadcast fragments for one specific root hash.
Lemma 3.1 ().
If honest nodes $v$ and $v^{\prime}$ broadcast fragments for root hashes $h$ and $h^{\prime}$, then $h=h^{\prime}$.
@@
Without loss of generality, let $h$ be the first hash for which $12t+$ 1 proposals are received at some node $v$. Thus, there are at least $t+1$ honest nodes that received their fragments from $sender$ and then broadcast their first proposal for root hash $h$.
Let $v^{\prime}$ be the first node that broadcasts a fragment for a hash $h^{\prime}\neq h$. Since it must hold that $|P(h^{\prime})|\geq 2t+1$ at node $v^{\prime}$, there must be at least $t+1$ honest nodes that broadcast the proposal for root hash $h^{\prime}$. Consequently, there must be an honest node $v^{*}$ that first proposed $h$ and then $h^{\prime}$. However, $v^{*}$ only sends a proposal for $h^{\prime}$ if $|F(h^{\prime})|\geq t+1$ according to Algorithm 2, which implies that it must have received a fragment for root hash $h^{\prime}$ from an honest node before. This is a contradiction to the assumption that $v^{\prime}$ is the first node that broadcasts such a fragment. ∎
@@
According to Algorithm 1, $v^{\prime}$ accepts messages associated with at most two different hashes for any sender $v$. After an initial proposal for some root hash $h$, $v$ may send a second proposal for $h^{\prime}$ if $|F(h^{\prime})|>=t+1$, i.e., there is at least one honest node that has broadcast a fragment for root hash $h^{\prime}$. Assume that $v$ sends another proposal for a different root hash $h^{\prime\prime}$, which again implies that at least one honest node must have broadcast a fragment for this root hash, a contradiction to Lemma 3.1.
Regarding the transmission of fragments, Lemma 3.1 also implies that an honest node only sends fragments for one root hash. Assume that $v$ sends fragments for a root hash $h^{\prime\prime}$ that differs from the hashes $h$ and $h^{\prime}$ for which it sends proposals. In this case, $v$ must have received at least $t+1$ fragments for one of the root hashes $h$ or $h^{\prime}$. Hence it follows that an honest node broadcast a fragment for a root hash other than $h^{\prime\prime}$, again contradicting Lemma 3.1. ∎
An important criterion for the totality property is that all honest eventually manage to get sufficiently many fragments of a message $m$ and proposals for the corresponding root hash if there is an honest node that delivers $m$. The last lemma states that this is indeed the case.
Lemma 3.3 ().
@@
Node $v$ adds all nodes from which it received fragments for root hash $h$ to the set $R(h)$. For any node $v^{\prime}\in R(h)$ it holds that $v^{\prime}$ either sent its own fragment or $v$’s fragment. However, the latter case also implies that $v^{\prime}$ must have its own fragment. According to Algorithm 1, $v$ sends fragment $f_{j}$ to $v_{j}$ for all $v_{j}\in V\setminus R(h)$, guaranteeing that the remaining nodes eventually receive and, due to Lemma 3.2, store their fragments as well. Thus, it eventually holds that $|P(h)|\geq 2t+1$ and $F(h)\neq\bot$ at all honest nodes, causing them to broadcast their fragments. Consequently, every honest node will eventually receive fragments of $m$ from at least $12t+$ 1 nodes. ∎
@@
Validity: If the sender $v$ of a message $m$ with root hash $h$ is honest, it sends the fragment $f_{j}$ to $v_{j}$ for all $v_{j}\in V$. Subsequently, every honest nodes broadcasts $proposal(h)$. Since all honest nodes eventually receive at least $12t+$ 1 proposals for hash $h$ (and only for hash $h$), they all broadcast their fragments. Thus, eventually $F(h)|\geq 2t+1$ and $|P(h)|\geq 2t+1$ holds at all honest nodes, triggering the delivery of $m$.
@@
It holds for algorithm $\mathcal{A}_{bit}$ that $\mathcal{L}(n,t)=2$.
@@
Let $|f|\coloneqq\frac{3}{2}|m|/n+O(\kappa\log(n))$ denote the size of the largest fragment message sent by an honest node. If the sender is honest, $|f|$ corresponds to the size of a fragment message containing a valid fragment of message $m$.
The sender may send fragment messages to all other nodes for a total of at most $(n-1)|f|$ bits. Each honest node may broadcast one fragment message containing its fragment and at most two proposals of size $O(\kappa)$. After reconstructing the message $m$, an honest node further sends a fragment message containing $f_{j}$ to all $v_{j}\in V\setminus R(h)$, where $|V\setminus R(h)|\leq t<n/3$. Not counting messages that nodes send to themselves, the communication complexity $\mathcal{C}(n,t,|m|)$ is therefore upper bounded by
@@
and therefore $\mathcal{L}(n,t)=\lim_{|m|\rightarrow\infty}\frac{1}{n|m|}\mathcal{C}(n,t,|m|)=2$. ∎
@@
If the sender is honest, all honest nodes receive their fragments and send the proposal for the corresponding root hash after at most $1$ time unit. Thus, after at most $2$ time units, all honest nodes get at least $12t+$ 1 proposals, triggering the broadcasting of fragments. After at most $3$ time units, the honest nodes receive the fragments from all honest nodes and deliver $m$. ∎
If the sender is Byzantine, an honest node may still deliver $m$ eventually. It is easy to see that it takes at most $3$ time units for every honest node to deliver after the first honest node delivers $m$.
@@
It is worth pointing out that the space complexity can be lowered to $\frac{3}{2}\ell_{max}+O(n\kappa)$ by introducing the rule that received fragments for different hashes from the same node are rejected. This rule is omitted from Algorithm 1 for ease of exposition.

### v2

@@ Abstract.
This paper introduces a novel mechanism to lower the communication and computational complexity. Two algorithms are presented that employ this mechanism to reliably broadcast messages in an asynchronous network where less than a third of all nodes are Byzantine. The first algorithm reduces the overhead factor to $2$ and has a time complexity of $3$ if the sender is honest, whereas the second algorithm attains an optimal time complexity of $2$ with the same overhead factor in the absence of equivocation. Moreover, an optimization is proposed that reduces the overhead factor to $3/2$ under normal operation in practice. Lastly, a lower bound is proved that an overhead factor lower than $3/2$ cannot be achieved for a relevant class of reliable broadcast algorithms.
@@ 1. Introduction
Curiously, there is little work on minimizing the constant in the $O(|m|n)$ term, although it likely dominates the actual bandwidth consumption in practice and is therefore crucial for real-world applications. The design by Cachin and Tessaro is based on a $(n,t+1)$-erasure code, where $t<n/3$ is the largest number of nodes that may exhibit Byzantine behavior. In this design, nodes broadcast encoded fragments of size $|m|/(t+1)\approx 3|m|/n$ instead of broadcasting $m$, sending at least $3|m|n$ bits altogether. An overhead factor of $3$ – compared to the ideal scenario where each node receives exactly $|m|$ bits – may already prove too costly for applications with strict bandwidth constraints. All schemes that follow this approach naturally inherit the overhead inherent in this design.
@@
Figure 1. Given an input of size 1 MiB, encoding and decoding times for erasure codes with parameters $(n,t+1)$ and $(n,2t+1)$ are shown for different network sizes. The circles and bars mark the averages and the 5th and 95th percentiles, respectively, of the measured times over 1000 runs.
As an added benefit, the transition from a $(n,t+1)$-erasure code to a $(n,2t+1)$-erasure code results in substantial performance improvements, both with respect to the encoding and decoding of fragments. Figure 1 shows performance numbers when encoding/decoding an input of size 1 MiB for various network sizes and the two different parameterizations. The experiments were conducted on an Intel Core i7 CPU with 6 cores and 32 GB of memory using an optimized library [^1] See https://github.com/AndersTrier/reed-solomon-simd. for Reed-Solomon codes (Reed and Solomon, 1960). The figure confirms the practicality of erasure coding due to its efficiency, exhibiting encoding and decoding times in the millisecond range, and scalability with respect to the network size. Moreover, the figure reveals that the encoding (decoding) time is lower by a factor of 2.2–2.4 (1.8–2.0) for all network sizes when using an $(n,2t+1)$-erasure code. Another interesting property is that the variance is also significantly lower, leading to more predictable encoding and decoding times.
The paper is organized as follows. The model is formally introduced in §2. A concrete algorithm that makes use of the novel mechanism is presented in §3. Its communication complexity is $2|m|n$ plus a term independent of $|m|$. Given an honest (i.e., non-faulty) sender, each honest node obtains the message after $3$ rounds of communication. A natural question is whether there is an algorithm with an overhead factor lower than $3$ and that terminates in an optimal $2$ rounds given an honest sender. In §4, this question is answered in the affirmative by presenting an algorithm with this property and the same overhead factor if the sender does not equivocate. Otherwise, the worst-case overhead factor is $\frac{5}{2}$ with respect to the largest fragment size transmitted by honest nodes. While the first algorithm only requires collision-resistant hash functions, the second algorithm makes use of threshold signatures. Both algorithms further bound the number of bits that are stored at each node in the worst case to a small multiple of a predetermined maximum message size. Practical considerations with the goal of reducing the communication complexity further in real-world deployments and simplifying the implementation are discussed in §5. A key observation is that the overhead can be reduced to $\frac{3}{2}$ during periods of synchrony and in the absence of failures. As shown in §5, a lower overhead factor than $\frac{3}{2}$ cannot be guaranteed for a particular class of algorithms. Related work is summarized in §6 before the paper concludes in §7.
@@ 2. Model
  • Validity: If the sender is honest and broadcasts $m$, then every honest node eventually delivers $m$.
@@
  • Totality: If an honest node delivers $m$, then all honest nodes eventually deliver $m$.
@@
In general, nodes may engage in multiple reliable broadcasts in parallel. In this case, the integrity condition effectively requires that messages are attributable to a uniquely identifiable execution context. In the following, we assume that every message implicitly contains an identifier and that each node runs a separate instance of the protocol for each identifier. The sender for a particular execution context is assumed to be globally known.
@@
As mentioned above, the objective is to reliably broadcast some message $m$, where the notation $|m|$ is used to denote the size of the message in bits. Instead of sending $m$, the algorithms introduced in §3 and §4 send fragments of $m$, which can be either chunks of $m$ itself or encoded data derived from $m$. Without loss of generality, we assume that there is some upper bound $\ell_{max}$ on the message size, which in turn limits the fragment size. State-of-the-art algorithms use either erasure codes, which can handle missing data, or error-correcting codes, which support the correction of erroneous data. The algorithms in this paper exclusively use an $(n,k)$-erasure codes with optimal reception efficiency, i.e., exactly $k$ out of $n$ symbols are sufficient to reconstruct a message of $k$ symbols. Throughout this paper, the parameters $n$ and $k$ correspond to the total number of nodes and the number of honest nodes, i.e., $k\coloneqq 2t+1$. We assume that each node has access to the routine get_fragments that takes a message $m$ as input and returns a list of $n$ fragments of size $\frac{|m|}{k}=\frac{|m|}{2t+1}<\frac{3}{2}\frac{|m|}{n}$ each. Furthermore, the nodes use the routine recover_message to recover $m$ given any subset $F$ of the $n$ fragments of size $|F|\geq k=2t+1$. This routine is assumed to always return some message, which may simply be a bit string of zeroes in case of an error, e.g., when the input contains fragments of different sizes.
@@
In addition to fragments, the nodes also send hashes, identifiers, and, in the case of the algorithm presented in §4, also signatures and signature shares. The hashes are assumed to be cryptographically strong in the sense that it is infeasible to find hash collisions (except with negligible probability). Naturally, it is also assumed that it is computationally infeasible to spoof signatures or signature shares. Since the cryptographic properties only hold for hashes, signatures, and shares of a certain minimum size, we introduce a security parameter $\kappa$ and define that the size of all data types other than fragments is bounded by $O(\kappa)$.
@@
Reliable broadcast algorithms are evaluated against multiple complexity measures. For a specific number $n$ of nodes and message size $\ell$, the communication complexity $\mathcal{C}(n,\ell)$ of an algorithm is the total number of bits sent by all honest nodes in the worst case assuming fewer than $n/3$ Byzantine nodes. If the sender is Byzantine, we define that $\ell$ is the message size corresponding to the largest fragment sent by an honest node. Given that at least $n\ell$ bits need to be transferred and we are concerned with the overhead for large messages, the primary goal is to minimize $\mathcal{L}(n)\coloneqq\lim_{\ell\rightarrow\infty}\frac{\mathcal{C}(n,\ell)}{\ell n}$. Additionally, the time complexity of an algorithm is of great practical importance, measuring the duration of an execution when normalizing the maximum message delay to $1$ time unit. According to Definition 2.1, an execution with a Byzantine sender may never terminate. Therefore, we restrict our attention to the case where the sender is honest. Due to the validity condition, this good-case time complexity must be bounded. Note that both presented algorithms ensure that the number of communication rounds is also bounded given a Byzantine sender. Lastly, the space complexity is considered as well, which is defined as the number of bits that any honest node stores during the execution of the algorithm in the worst case.
@@ 3.2. Description
Each node executing $\mathcal{A}_{bit}$ locally maintains the data structures $F$, $R$, $H$, and $P$. Let $\mathcal{F}$, $\Pi$, and $\mathcal{H}$ denote the set of all possible fragments, Merkle proofs, and hashes, respectively. The hash map $F:\mathcal{H}\times V\rightarrow\mathcal{F}\times\Pi$ stores the fragment $f_{j}\in\mathcal{F}$ and Merkle proof $\pi_{j}\in\Pi$ of node $v_{j}$ for the message with the root hash $h\in\mathcal{H}$. If $(f_{j},\pi_{i})$ is locally available, then $F(h,v_{j})=(f_{j},\pi_{j})$, and $F(h,v_{j})=\bot$ otherwise. This hash map is used to collect fragments with the goal of eventually recovering the corresponding message. We further define that $F(h)\coloneqq\cup_{v_{j}\in V}F(h,v_{j})$ denotes the set of all collected $(v_{j},\pi_{j})$ pairs for hash $h$. The hash map $R:\mathcal{H}\rightarrow 2^{V}$ stores the nodes from which a fragment for root hash $h$ has been obtained. This hash map is needed to determine which nodes may still be missing their fragments. Conversely, for any node $v$, the hash map $H:V\rightarrow 2^{\mathcal{H}}$ contains the set of hashes for which a proposal or fragment was received from node $v$. The purpose of this data structure is to bound both the communication and space complexity. Lastly, $P:\mathcal{H}\rightarrow 2^{V}$ indicates which nodes have proposed the delivery of the message associated with the Merkle root hash $h$. The collected proposals are used to ensure that only fragments with sufficient support are broadcast. All hash maps are initially empty. In addition to the hash maps, each node further uses the Boolean variable $done$, initially $false$, to capture the information whether the execution has terminated, either with or without the delivery of a message. In the latter case, the totality condition implies that no honest node will deliver a message for this execution. Moreover, let $\mathbf{H}\coloneqq\cup_{v\in V}H(v)$ and $h_{max}\coloneqq\arg\max_{h\in\mathbf{H}}|P(h)|$, breaking ties arbitrarily if there is no single hash for which the most proposals have been received.
@@
Algorithm 1 $\mathcal{A}_{bit}$: Triggered actions at node $v_{i}$. Initially, $F=R=H=P=\{\}$.
@@
Algorithm 2 describes the state-based actions of algorithm $\mathcal{A}_{bit}$. An honest node $v_{i}$ broadcasts its fragment $f_{i}$, at most once, after collecting $12t+$ 1 proposals for the corresponding Merkle root hash $h$. By contrast, a node $v_{i}$ broadcasts a proposal for a specific root hash $h$ not only when it receives its fragment from the $sender$ but also when at least $t+1$ fragments for the root hash $h=h_{max}$ have been received and $v_{i}$ has not broadcast the proposal before. Lastly, if a node $v_{i}$ receives at least $12t+$ 1 fragments and proposals for $h_{max}$, it recovers the message $m$ and, additionally, recomputes the fragments and the corresponding root hash. If the computed hash matches $h_{max}$, node $v_{i}$ concludes that the fragments are valid and the message $m$ can be delivered. In this case, $v_{i}$ first recomputes the Merkle proof $\pi_{j}$ and sends $fragment(h_{max},j,f_{j},\pi_{j})$ to each $v_{j}\in V\setminus R(h_{max})$ before calling deliver($m$) to deliver the message. Note that $done$ is set to $true$ after executing these steps even if $h\neq h_{max}$ because it is computationally infeasible to find any $12t+$ 1 fragments such that equality holds.
@@
Algorithm 2 $\mathcal{A}_{bit}$: State-based actions at node $v_{i}$. Initially, $F=R=P=\{\}$, $done=false$. Let $h_{max}\coloneqq\arg\max_{h\in\mathbf{H}}|P(h)|$.
@@ 3.3. Analysis
In this section, we prove the correctness of $\mathcal{A}_{bit}$ as well as its communication, time, and space complexity. A series of lemmas is used to simplify the proof structure. The first lemma is concerned with the fragments that honest nodes broadcast, showing that the proposal mechanism ensures that honest nodes only broadcast fragments for one specific root hash.
Lemma 3.1.
If honest nodes broadcast fragments for root hashes $h$ and $h^{\prime}$, then $h=h^{\prime}$.
@@
Without loss of generality, let $h$ be the first hash for which $12t+$ 1 proposals are received at some node $v$. Thus, there are at least $t+1$ honest nodes that received their fragments from the $sender$ and then broadcast their first proposal for root hash $h$.
Let $v^{\prime}$ be the first node that broadcasts a fragment for a hash $h^{\prime}\neq h$. Since it must hold that $|P(h^{\prime})|\geq 2t+1$ at node $v^{\prime}$, there must be at least $t+1$ honest nodes that broadcast the proposal for root hash $h^{\prime}$. Consequently, there must be an honest node $v^{*}$ that first proposed $h$ and then $h^{\prime}$. However, $v^{*}$ only sends a proposal for $h^{\prime}$ if $|F(h^{\prime})|\geq t+1$ according to Algorithm 2, which implies that an honest node must have sent a fragment for root hash $h^{\prime}$ before, a contradiction to the assumption that $v^{\prime}$ is the first such node. ∎
@@
According to Algorithm 1, $v^{\prime}$ accepts messages associated with at most two different hashes for any sender $v$. After an initial proposal for some root hash $h$, $v$ may send a second proposal for $h^{\prime}$ if $|F(h^{\prime})|\geq t+1$, i.e., there is at least one honest node that has broadcast a fragment for root hash $h^{\prime}$. Assume that $v$ sends another proposal for a different root hash $h^{\prime\prime}$, which again implies that at least one honest node must have broadcast a fragment for this root hash, a contradiction to Lemma 3.1.
Regarding the transmission of fragments, Lemma 3.1 also implies that an honest node only sends fragments for one root hash. Assume that $v$ sends fragments for a root hash $h^{\prime\prime}$ that differs from the hashes $h$ and $h^{\prime}$ for which it sends proposals. In this case, $v$ must have received at least $t+1$ fragments for root hash $h$ or $h^{\prime}$. Hence it follows that an honest node broadcast a fragment for a root hash other than $h^{\prime\prime}$, again contradicting Lemma 3.1. ∎
An important criterion for the totality property is that all honest eventually obtain sufficiently many fragments of a message $m$ and proposals for the corresponding root hash if there is an honest node that delivers $m$. The last lemma states that this is indeed the case.
Lemma 3.3.
@@
Node $v$ adds all nodes from which it received fragments for root hash $h$ to the set $R(h)$. For any node $v^{\prime}\in R(h)$ it holds that $v^{\prime}$ either sent its own fragment or $v$’s fragment. However, the latter case also implies that $v^{\prime}$ must have its own fragment. According to Algorithm 1, $v$ sends $f_{j}$ to $v_{j}$ for all $v_{j}\in V\setminus R(h)$, guaranteeing that these nodes eventually receive and, due to Lemma 3.2, store their fragments as well. Thus, it eventually holds that $|P(h)|\geq 2t+1$ and $F(h)\neq\bot$ at all honest nodes, causing them to broadcast their fragments. Consequently, every honest node will eventually receive at least $12t+$ 1 fragments of $m$. ∎
@@
Validity: If the sender $v$ of a message $m$ with root hash $h$ is honest, it sends the fragment $f_{j}$ to $v_{j}$ for all $v_{j}\in V$. Subsequently, every honest nodes broadcasts $proposal(h)$. Since all honest nodes eventually receive at least $12t+$ 1 proposals for hash $h$ (and only for hash $h$), they all broadcast their fragments. Thus, eventually $|F(h)|\geq 2t+1$ and $|P(h)|\geq 2t+1$ holds at all honest nodes, triggering the delivery of $m$.
@@
It holds for algorithm $\mathcal{A}_{bit}$ that $\mathcal{L}(n)=2$.
@@
Let $|f|\coloneqq\frac{3}{2}|m|/n+O(\kappa\log(n))$ denote the size of the largest fragment message sent by an honest node. If the sender is honest, $|f|$ corresponds to the size of a fragment message containing a valid fragment of message $m$. The sender may send fragment messages to all other nodes for a total of at most $(n-1)|f|$ bits. Each honest node may broadcast one fragment message containing its fragment and at most two proposals of size $O(\kappa)$. After reconstructing the message $m$, an honest node further sends a fragment message containing $f_{j}$ to all $v_{j}\in V\setminus R(h)$, where $|V\setminus R(h)|\leq t<n/3$. Not counting messages that nodes send to themselves, the communication complexity $\mathcal{C}(n,|m|)$ is therefore upper bounded by
@@
and therefore $\mathcal{L}(n)=\lim_{|m|\rightarrow\infty}\frac{1}{n|m|}\mathcal{C}(n,|m|)=2$. ∎
@@
If the sender is honest, all honest nodes receive their fragments and send the proposal for the corresponding root hash after at most $1$ time unit. Thus, after at most $2$ time units, all honest nodes get at least $12t+$ 1 proposals and broadcast their fragments. After at most $3$ time units, the honest nodes receive at least $12t+$ 1 fragments and deliver the message. ∎
If the sender is Byzantine, an honest node may still deliver $m$ eventually. It is easy to see that every honest node delivers $m$ at most $3$ time units later in this case.
@@
Note that the space complexity can trivially be lowered to $\frac{3}{2}\ell_{max}+O(n\kappa)$ by introducing the rule that received fragments for different hashes from the same node are rejected. This rule is omitted from Algorithm 1 for ease of exposition.